OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4).

OpenStudy (zzr0ck3r):

form is y=mx+b m = slope so y=-3x+b use the given point to find b -4=-3(3)+b -4=-9+b b = 5 y=-3x+5